# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ 4 Network for AI
The current network has achieved preliminary success in advancing AI4NET, developing a series of AI use cases and scenarios, and partially implementing standardization efforts. However, most of these AI4NET implementations adopt an “add-on” design pattern, which, while convenient, may introduce additional challenges (such as latency, data privacy concerns, etc.), hindering the network’s evolution towards more efficient AI services to a certain extent. The underlying reason lies in the insufficient support the current network architecture provides for deeply integrating AI into the network. Although 5G exhibits a trend towards cloud-native, IT software-defined Service virtualization (exemplified by the adoption of Service-Based Architecture Core Network and cloud-native deployment of core networks), 3GPP has standardized the Network Data Analytics Function (NWDAF), Management Data Analytics Function, and related mechanisms and processes, with attempts to define and introduce standardized AI operational paradigms for critical use cases in RAN3 and RAN1, limitations persist due to factors such as technical maturity, security concerns, and the complexity of system operation and maintenance. As a result, the access network side largely retains the conventional “siloed CT-based BS architecture” and a relatively rigid protocol stack model. From an end-to-end perspective of AI service flows, there is a lack of defined AI functional components on the access network side, and the AI service functionalities between the core network and access network have yet to be integrated or aligned. This overall design remains incomplete, with vulnerabilities and deficiencies hindering the current network architecture from supporting the intelligent vision of 6G.
To support “native intelligence” and achieve ubiquitous general intelligence, the 6G network necessitates groundbreaking innovative designs at the network architecture level, enabling the endogenous embedding of AI capabilities within the network. This represents the core concept of Phase II (i.e., NET4AI) in the deep integration of 6G and AI, signifying a shift from a communication network solely providing connectivity services to a new network paradigm that offers a diverse array of services encompassing connectivity, computing power, data, and algorithms. Through this new paradigm, encompassing core networks, access networks, and terminals, the network leverages resources and functionalities, such as connectivity, computation, data, and models, to bolster AI with multiple support capabilities. This enables AI training/inference to be executed more efficiently in real time or enhances data security and privacy protection, among other benefits.
@@ 4.1 NET4AI Architecture over 6G
China Mobile has proposed a systematic 6G network architecture design that transcends connectivity in the industry, centered on the core concept of “three layers and six planes” [150]. The “three layers” encompass the Physical Resource Layer, the Network Function Layer, and the Application & Service Layer. At the same time, the “six planes” comprise the Communication Plane, the Data Plane, the Computing Plane, the Intelligence Plane, and the Security Plane. This architecture focuses on multi-dimensional capability elements within the network, aiming to provide comprehensive services throughout the information flow process. China Telecom has proposed a ubiquitous and ultra-converge 6G network architecture, built upon the core concept of “three layers and three sectors” [151]. The “three layers” consist of the Infrastructure Layer, the Network Function Layer, and the Network Operating System Layer. The “three sectors” embody Data Integration, Intelligent Simplicity, and Trustworthiness. This architecture leverages the Intelligent Simplicity sector to integrate the intelligent brain, connections, and facilities across the three layers, thereby enabling intelligent inclusiveness. China Unicom’s vision for future 6G networks is summarized as intelligent, converged, green, and trustworthy, structured vertically into three vertical communication layers of land-based, air-based, and space-based, as well as three horizontal domains of Perception Resource Domain, Function Control Domain, and Service Application Domain [152]. This structure is interconnected through two chains: Intelligent Native and Secure & Trustworthy.
@@
Huawei has unveiled a task-centric AI architecture of 6G networks [153], asserting that 6G networks will introduce new resource dimensions and support the coordination of multi-dimensional resources, enabling 6G networks to natively support AI and achieve a transformation from session-centric to task-centric operations. CICT puts forward the concept of multi-network integration and ubiquitous intelligence, believing that 6G would realize global three-dimensional deep coverage, develop body area networks centered on human beings, form multi-layer coverage with both breadth and depth and learn the ubiquitous intelligence system with multi-network integration [154, 155]. ZTE has identified intelligent radio, coverage, and evolution as the fundamental technical characteristics of 6G network architecture. Enabling technologies for 6G will include three-dimensional connectivity, intelligent MIMO, on-demand topology, on-demand AI, and new horizon communications. OPPO indicates that 6G will reshape human interaction and AI, making AI a ubiquitous infrastructure. The 6G system architecture will be coordinated by intelligence, performance, and flexibility, deeply integrating AI capabilities into the user plane, control plane, and function plane [156]. The IMT-2030 (6G) Promotion Group has envisioned the 6G network architecture as an open and innovative platform for information services [157], offering capabilities that transcend mere connectivity. These capabilities encompass computing power networks, trusted security, sensing, data services, and AI-based network intelligent autonomy.
@@
Ericsson is focused on researching cognitive networks and leveraging AI to achieve data-driven security and highly automated network operations [158]. Qualcomm advocates that AI and ML will fundamentally transform the design and deployment of wireless communication and networking systems. As the leader of the EU’s 6G project Hexa-X, Nokia states that AI will be the primary driving force behind the technological transformation of the new radio interface for 6G. The company is creating unique AI use cases and scenarios and developing foundational AI technologies [159]. LG Electronics is focused on the research and development of 6G AI technology, anticipating that 6G systems will usher in the era of the “Internet of Everything and Environment” powered by AI. Japan’s “Beyond 5G Promotion Strategy” project aims to introduce AI into network management, enabling network autonomy and driving network evolution towards intelligence.
@@
Zhang et. al. at Beijing University of Posts and Telecommunications (BUPT), put forth a 6G network design philosophy of “Human-Machine-Thing-Spirit”, emphasizing the inevitable trend of AI integration in 6G, leveraging cognitive enhancement and decision-making simulation to intelligently define network requirements, ensuring secure and reliable network transmission, and realizing intent-driven networks [160]. Cui et. al. also from BUPT, designed a fusion of 6G RAN and AI, summarizing its four key characteristics as green, multi-dimensional, stereo, and full-scenario service [161]. You et. al. at the Purple Mountain Laboratories and Southeast University proposed that 6G will integrate communication, computation, control/caching, and AI to enable full-application scenarios, with 6G networks being inherently secure [162]. Liang et. al. at the University of Electronic Science and Technology of China (UESTC) demonstrated the vision of integrating 6G and AI and studied implementation examples in wireless communications based on two classic AI algorithms: deep learning and reinforcement learning [163]. Zhang et. al. at Zhejiang University (ZJU) and Zhejiang Lab explained that AI will become the cornerstone of 6G, and “Intelligent inclusion” will become an essential feature of 6G [164]. The 6G network will be able to provide intelligent services to all types of end users. Tao et. al. at Shanghai Jiaotong University (SJTU) indicated that the typical technical feature of 6G is the intelligence of the network, which is not only reflected in the network itself being able to self-learn and self-optimize but also being able to provide intelligent services to the outside world [165]. Among them, 6G edge intelligence is expected to become an important research topic. Cui et. al. at the Chinese University of Hong Kong (Shenzhen) and Shenzhen Research Institute of Big Data pointed out that in the 6G era, networks and AI will be fully integrated, and customized AI services will be provided on the wireless network side [166, 167]. Gao et. al. at Tsinghua University (THU) advocated that AI will be used in wireless communications and change the top-level architecture of wireless networks [168]. At the University of Oslo in Norway, Zhang Yan and his team reported that 6G would enhance its progress in intelligence, especially in edge intelligence [169]. Among them, integrating DT and edge networks will be a promising attempt. At the University of Oulu in Finland, Taleb and his team initiated the 6G Flagship research program, advocating for realizing context-aware intelligent services and applications for human and non-human users through AI [170].
@@
Incorporating the aforementioned research ideas and architectural visions, the industry has reached a core understanding of the intelligent design of 6G network architectures, emphasizing “distributed structure”, “endogenous intelligence”, and “integrated simplicity”. Based on this consensus, future networks will evolve into advanced, integrated platforms that surpass mere communication connectivity services, offering multi-dimensional services and corresponding network capabilities beyond communication. Generally speaking, the consensus suggests that 6G networks are expected to provide communication connectivity, data and services, computing and services, security and services, AI and services, and management and orchestration capabilities and services. Furthermore, they will cater to diverse businesses by offering end-to-end lifecycle management services throughout their entire process. These consensuses are illustrated in Figure 15.
@@
Communication Connectivity Capability: In 6G, the Control Plane (CP) and User Plane (UP) within the communication layer will undergo further enhancement designs to deepen the service orientation and openness of network functions to address the complex and diverse business scenarios and emerging demands of the future.
(1) Enhanced CP: Within the CP, 6G necessitates enhancing and optimizing the service management and interfaces of existing CP functions to reduce the complexity of network function and signaling interaction processes. Additionally, 6G will extend the traditional CP by integrating space-terrestrial ubiquitous access control management, enabling new functionalities and service-oriented offerings for session management and policy control across communication, computing, data, and other essential elements.
(2) Enhanced UP: The enhancement of the UP focuses on augmenting or introducing new, more generalized, and flexible UP forwarding protocols that can perform agile service forwarding and processing based on communication, computing, and other elements. Leveraging microservice technologies and next-generation hardware support, e.g., high-performance programmable chips, cloud-native technologies, and the P4 (Programming Protocol-Independent Packet Processors) advanced language, UP functions will be restructured into service-oriented modules. This flexible invocation and efficient collaboration among functional modules will break the constraints of traditional layered protocols, enhancing data forwarding efficiency and bandwidth utilization.
Data Capability: The traditional UP architecture struggles to support distributed management and efficient utilization in response to the diverse and highly time-sensitive requirements of sensory data, AI data, network behavior, status data, and other forms. Consequently, 6G networks necessitate a novel data capability that differs from traditional UPs. This capability must adhere to data regulation and supervision requirements while enhancing data analysis and processing efficiency, addressing data management challenges, and enabling value realization.
@@
Security Capability: Traditional security measures have struggled to meet the complex and ever-changing security demands of 6G. Currently, the design of security mechanisms stands independently from the network architecture, leaving room for improvement in both security and efficiency. Hence, 6G networks necessitate establishing an inherently trustworthy and self-driving security system at the core of the system. The rapid advancements in cloud computing, big data, and AI technologies offer technical support for constructing a 6G network security system. The 6G security system will be structured around three key components: security capabilities, security controls, and security decision-making. The security capability layer deploys security resources (including physical security resources and pooled security capabilities) on demand, enabling flexible scheduling. The security control layer oversees the scheduling of security resources and orchestrates trust and security collaborations. Meanwhile, the security decision-making layer integrates AI analytics capabilities with DT networks to facilitate intelligent security analysis, simulation, and the integration and delivery of security service capabilities. Based on this architecture, 6G networks will foster a trustworthy and inherently secure system characterized by intelligent proactive defense, ubiquitous intelligent collaboration, and robust privacy protection.
@@
AI Capability: Aligned with the trend of network intelligence evolution, the 6G network architecture will support ubiquitous and inclusive AI capabilities. This necessitates a breakthrough from the external or patch-based intelligence of 5G, transforming intelligence into an inherent network ability. Consequently, the AI capabilities in 6G networks refer to those that cater to the network’s operation and maintenance and provide AI services to users or externally. By endogenously constructing AI elements (e.g., computing power, algorithms, data, etc.) and under centralized intelligent orchestration, management, control, and distributed intelligent execution, 6G enables the on-demand customization of AI services.
The network side will provide lifecycle management functions for AI models and data, encompassing capabilities such as AI data management (e.g., training data reporting, gradient reporting), AI model management (e.g., model generation, storage, selection, distribution, updating, replacement, activation/deactivation, and performance monitoring), and AI computing power allocation (e.g., for training and inference execution).
@@
Management &\&& Orchestration Capability: The immense logical structure and technological ecosystem complicate network operation and maintenance management, necessitating high network autonomy in 6G networks. Consequently, there is an urgent need for management and orchestration capabilities to uniformly and dynamically orchestrate and schedule various functional components based on business requirements, enabling self-optimization, self-evolution, and self-healing of the network. The 6G system will introduce external applications and service demands while exposing internal diversified resources and functions. This requires flexible cross-functional flow and scheduling of multi-dimensional resources, such as computing, communication, and storage within the network, offering customized and personalized application services based on virtualization and microservice technologies. For instance, AI capabilities and analyzed data within the network can be opened to third parties to provide services and necessary support, a crowdsourcing behavior that necessitates the support of management and orchestration functions. Furthermore, through joint analysis and prediction of computing resources and network traffic, containerized gNBs can automatically scale up or down based on predicted network traffic loads, enabling dynamic resource allocation and network energy savings. Therefore, management and orchestration capabilities are closely related to all other capabilities, orchestrating and managing resources and functions efficiently and flexibly across all functional layers of the 6G network, achieving joint optimization.
@@
Within the consensus of the 6G network architecture, communication connectivity is responsible for establishing end-to-end communication links and providing precise and highly reliable connectivity services. The data capability oversees data collection, cleansing, processing, and storage within the end-to-end network, offering data subscription and update services to other planes. The computation capability is tasked with scheduling and optimizing various computational resources within the network to handle computational tasks. The security capability provides inherent security support for the entire network, offering security services to customers while opening up security capabilities to external partners or facilitating crowdsourcing of external security capabilities. The AI capability provides an AI engine that, in conjunction with functional requirements across various domains within the network, offers big data analytics, AI algorithms, model training services, and simulation verification of related solutions. AI capabilities can be centralized or distributed in network elements and terminals to support real-time AI applications or in the cloud, leveraging vast cloud data for complex algorithm implementations. Additionally, the network’s AI capabilities can be opened to the outside world, while external AI capabilities can be introduced into the network, enabling the crowdsourcing of AI capabilities. The management & orchestration capability is responsible for scheduling various capabilities within the network to achieve flexible organization, fulfilling the orchestration and management of different resources and functions to meet users’ personalized service demands.
@@
Figure 16: Architecture diagram of NET4AI. The 6G network physically comprises three entity networks: the cloud core network (including centralized data/computation centers), the edge-side access network (including 3GPP and non-3GPP connection methods), and the terminal-side access network (including wired and wireless connection methods), as depicted in the lower half of the figure. NET4AI represents the mapping and supporting role of communication connectivity, data, computing, security, and management orchestration within AI, among the six capabilities and services (i.e., communication connectivity, data, computation, security, AI, and management & orchestration) abstracted from the 6G network’s functional services, as indicated by the arrows in the upper half of the figure.

### v3

@@ 3 Network for AI
At present, preliminary achievements have been made in promoting AI4NET. A series of AI use cases and scenarios have been developed, and part of the standardization work has been accomplished under the leadership of 3GPP. This has led to the formation of various module-level “plug-in” AI functions in 5G networks, which are used to optimize the efficiency of wireless transmission and improve system performance for identified and specific communication problems. Although this approach is convenient, it still lacks considerations of systematicity and interpretability, which greatly limits the scalability of AI capability and the ability of model generalization, and hinders the development of more efficient AI services in the networks. Essentially, this is because the current network architecture cannot provide sufficient support for the in-depth integration of AI into the network.
Although key technologies such as cloud-native, information technology (IT) software-defined, and service virtualization have been developed during the evolution of 5G networks (for example, CN that supports service-based architecture and cloud-based deployment of CN functions, etc.), and 3GPP has standardized and defined the Network Data Analytics Function (NWDAF) as AI functional components in the CN, along with relevant mechanism processes, and is attempting to introduce and define standardized AI operation paradigms. However, these efforts are still restricted by multiple conditions including technology maturity, security, and the complexity of system O&M. As a result, the current access network architecture generally still retains the traditional “siloed communication-technicalized BS architecture”, i.e., the current BS architecture is still highly integrated and designed based on communication services, and cannot effectively connect and adapt to emerging AI services. This leads to the imperfection of the overall network design, with relatively major loopholes and deficiencies remaining. Because from the perspective of the end-to-end AI service process, there is no corresponding definition of AI functional components on the access network side, and the docking and integration of AI service functions between CN and the access network have not been achieved, resulting in the current network architecture being unable to support the intelligent vision of 6G.
To support “native intelligence” and achieve ubiquitous general intelligence, the 6G network requires groundbreaking innovative innovations in the network architecture to enable the endogenous embedding of AI capabilities within the network. This represents the core concept of Phase II (i.e., NET4AI) in the deep integration of 6G and AI, signifying a shift from a communication network solely providing connectivity services to a new network paradigm that offers a diverse array of services, including connectivity, computing power, data, and algorithms. This new paradigm would span CNs, access networks, and terminals, to provide comprehensive support of AI. This enables AI training/inference to be executed more efficiently in real time while enhancing data security and privacy protection, among other benefits.
@@ 3.1 NET4AI Architecture over 6G
China Mobile in [110] proposed a systematic 6G network architecture design that transcends connectivity in the industry, centered on the core concept of “three layers and six planes”. The “three layers” refers to the physical resource layer, the network function layer, and the application & service layer, while the “six planes” comprise the communication plane, the data plane, the computing plane, the intelligence plane, the security plane, and the management&orchestration plane. The first five planes are independent functional planes formed by the decoupling of the 6G network from end to end, providing corresponding capabilities and services. The management&orchestration plane realizes the flexible combination of functions across layers and planes, realizes the management and scheduling of multiple resources, and forms a completes service chain. This architecture focuses on multi-dimensional capability elements within the network, aiming to provide comprehensive services throughout the information flow process. In [111], China Telecom proposed a ubiquitous and ultra-converge 6G network architecture, built upon the core concept of “three layers and three sectors”. The “three layers” consist of the Infrastructure Layer, the Network Function Layer, and the Network Operating System Layer, while the “three sectors” embody Data Integration, Intelligent Simplicity, and Trustworthiness. This architecture leverages the Intelligent Simplicity sector to integrate the intelligent brain, connections, and facilities across the three layers, thereby enabling intelligent inclusiveness. China Unicom in [112] envisioned furture 6G network as intelligent, converged, green, and trustworthy. Its architecture is structured vertically into three communication layers of land-based, air-based, and space-based and horizontally into three domains of Perception Resource Domain, Function Control Domain, and Service Application Domain. These layers and domains are interconnected by two chains: Intelligent Native and Secure & Trustworthy.
@@
Huawei in [113] unveiled a task-centric AI architecture of 6G networks, asserting that 6G networks will introduce new resource dimensions and support the coordination of multi-dimensional resources. This design allows 6G networks to natively support AI and achieve a transformation from session-centric to task-centric operations. CICT in [114, 115] put forward the concept of multi-network integration and ubiquitous intelligence, envisioning that 6G will realize global three-dimensional deep coverage, develop body area networks centered on human beings, form multi-layer coverage with both breadth and depth and learn the ubiquitous intelligence system with multi-network integration. OPPO in [116] suggested that 6G will redefine human interaction and AI, by making AI a ubiquitous infrastructure. The 6G system architecture will emphasize the coordination across intelligence, performance, and flexibility, and integrate deeply AI capabilities into the user plane, control plane, and function plane. More authoritative and official, The IMT-2030 (6G) Promotion Group in China envisioned the 6G network architecture in [117], as an open and innovative platform for information services, offering capabilities that transcend mere connectivity. These capabilities encompass computing power networks, trusted security, sensing, data services, and AI-based network intelligent autonomy.
@@
Ericsson in [118] is focused on researching cognitive networks and leveraging AI to achieve data-driven security and highly automated network operations. Qualcomm advocated that AI and ML will fundamentally transform the design and deployment of wireless communication and networking systems. As the leader of the EU’s 6G project Hexa-X, Nokia in [119] stated that AI will be the primary driving force behind the technological transformation of the NR interface for 6G. The company is creating unique AI use cases and scenarios and developing foundational AI technologies. LG Electronics is focused on the research and development of 6G AI technology, anticipating that 6G systems will usher in the era of the “Internet of Everything and Environment” powered by AI. Japan’s “Beyond 5G Promotion Strategy” project aimed to introduce AI into network management, enabling network autonomy and driving network evolution towards intelligence.
@@
Following the industrial perspectives on AI integrated 6G network architecture, academic research has also proposed various innovative ideas for shaping the future 6G network architecture. The authors put forth a 6G network design philosophy of “Human-Machine-Thing-Spirit”[120], emphasizing the inevitable trend of AI integration in 6G, leveraging cognitive enhancement and decision-making simulation to intelligently define network requirements, ensuring secure and reliable network transmission, and realizing intent-driven networks. The authors of [121] introduced a fusion of 6G RAN and AI, identifying four key characteristics as green, multi-dimensional, stereo, and full-scenario service. The authors proposed that 6G will integrate communication, computation, control/caching, and AI to support full-application scenarios, and will be inherently secure[122]. The authors of [123] demonstrated a vision of integrating 6G and AI and studied implementation examples in wireless communications based on two classic AI algorithms: DL and RL. In [30], the authors argued that AI will become the cornerstone of 6G with “Intelligent inclusion” serving as an essential feature of 6G, enabling the network to provide intelligent services to all types of end users. The authors of [124] advocated that AI will be used in wireless communications and change the top-level architecture of wireless networks. The authors of [125] emphasized that the NET4AI paradigm should leverage the capabilities of 6G networks, such as precise wireless sensing, high data-rate transmission, and ubiquitous connectivity, for intelligence distillation at the network edge, promoting edge AI and, therefore enabling ubiquitous and trustworthy AI applications in wireless networks. Among them, federated edge learning (FEEL) is expected to become an important learning architecture over 6G. In [126, 127], the authors pointed out that in the 6G era, networks and AI will be fully integrated, and customized AI services will be provided on the wireless network side. The authors of [128] reported that 6G would enhance its progress in intelligence, especially in edge intelligence. Among them, integrating DT and edge networks will be a promising attempt. The authors of [129] initiated the 6G Flagship research program, advocating for realizing context-aware intelligent services and applications for human and non-human users through AI.
@@
Incorporating the aforementioned research ideas and architectural visions, the industry has reached a core understanding of the intelligent design of 6G network architectures, emphasizing “distributed structure”, “endogenous intelligence”, and “integrated simplicity”. Based on this consensus, future networks will evolve into advanced, integrated platforms that surpass mere communication connectivity services, offering multi-dimensional services and corresponding network capabilities beyond communication. Generally speaking, the consensus suggests that 6G networks are expected to provide communication connectivity, data capability, computing capability, security capability, AI capability, management and orchestration capability and their corrsponding services. Furthermore, they will cater to diverse businesses by offering end-to-end lifecycle management services throughout their entire process. These consensuses are illustrated in Figure 13.
@@
Communication Connectivity Capability: In 6G, the control plane and user plane within the communication layer will undergo further enhancement designs to deepen the service orientation and openness of network functions to address the complex and diverse business scenarios and emerging demands of the future.
(1) Enhanced control plane: Within the control plane, 6G necessitates enhancing and optimizing the service management and interfaces of existing control plane functions to reduce the complexity of network function and signaling interaction processes. Additionally, 6G will extend the conventional control plane by integrating space-terrestrial ubiquitous access control management, enabling new functionalities and service-oriented offerings for session management and policy control across communication, computing, data, and other essential elements.
(2) Enhanced user plane: The enhancement of the user plane focuses on augmenting or introducing new, more generalized, and flexible user plane forwarding protocols that can perform agile service forwarding and processing based on communication, computing, and other elements. Leveraging microservice technologies and next-generation hardware support, e.g., high-performance programmable chips, cloud-native technologies, and the P4 (programming protocol-independent packet processors) advanced language, user plane functions (UPF) will be restructured into service-oriented modules. This flexible invocation and efficient collaboration among functional modules will break the constraints of conventional layered protocols, enhancing data forwarding efficiency and bandwidth utilization.
Data Capability: The conventional user plane architecture struggles to support distributed management and efficient utilization in response to the diverse and highly time-sensitive requirements of sensory data, AI data, network behavior, status data, and other forms. Consequently, 6G networks necessitate a novel data capability that differs from conventional user plane. This capability must adhere to data regulation and supervision requirements while enhancing data analysis and processing efficiency, addressing data management challenges, and enabling value realization.
@@
Security Capability: conventional security measures have struggled to meet the complex and ever-changing security demands of 6G. Currently, the design of security mechanisms stands independently from the network architecture, leaving room for improvement in both security and efficiency. Hence, 6G networks necessitate establishing an inherently trustworthy and self-driving security system at the core of the system. The rapid advancements in cloud computing, big data, and AI technologies offer technical support for constructing a 6G network security system. The 6G security system will be structured around three key components: security capabilities, security controls, and security decision-making. The security capability layer deploys security resources (including physical security resources and pooled security capabilities) on demand, enabling flexible scheduling. The security control layer oversees the scheduling of security resources and orchestrates trust and security collaborations. Meanwhile, the security decision-making layer integrates AI analytics capabilities with DT networks to facilitate intelligent security analysis, simulation, and the integration and delivery of security service capabilities. Based on this architecture, 6G networks will foster a trustworthy and inherently secure system characterized by intelligent proactive defense, ubiquitous intelligent collaboration, and robust privacy protection.
@@
AI Capability: Aligned with the trend of network intelligence evolution, the 6G network architecture will support ubiquitous and inclusive AI capabilities. This necessitates a breakthrough from the external or patch-based intelligence of 5G, transforming intelligence into an inherent network ability. Consequently, the AI capabilities in 6G networks refer to those that cater to the network’s O&M and provide AI services to users or externally. By endogenously constructing AI elements (e.g., computing power, algorithms, data, etc.) and under centralized intelligent orchestration, management, control, and distributed intelligent execution, 6G enables the on-demand customization of AI services. The network side will provide lifecycle management functions for AI models and data, encompassing capabilities such as AI data management (e.g., training data reporting, gradient reporting), AI model management (e.g., model generation, storage, selection, distribution, updating, replacement, activation/deactivation, and performance monitoring), and AI computing power allocation (e.g., for training and inference execution).
@@
Management &\&& Orchestration Capability: The immense logical structure and technological ecosystem complicate network O&M management, necessitating high network autonomy in 6G networks. Consequently, there is an urgent need for management and orchestration capabilities to uniformly and dynamically orchestrate and schedule various functional components based on business requirements, enabling self-optimization, self-evolution, and self-healing of the network. The 6G system will introduce external applications and service demands while exposing internal diversified resources and functions. This requires flexible cross-functional flow and scheduling of multi-dimensional resources, such as computing, communication, and storage within the network, offering customized and personalized application services based on virtualization and microservice technologies. For instance, AI capabilities and analyzed data within the network can be opened to third parties to provide services and necessary support, a crowdsourcing behavior that necessitates the support of management and orchestration functions. Furthermore, through joint analysis and prediction of computing resources and network traffic, containerized gNBs can automatically scale up or down based on predicted network traffic loads, enabling dynamic resource allocation and network energy savings. Therefore, management and orchestration capabilities are closely related to all other capabilities, orchestrating and managing resources and functions efficiently and flexibly across all functional layers of the 6G network, achieving joint optimization.
@@
Within the consensus of the 6G network architecture, communication connectivity is responsible for establishing end-to-end communication links and providing precise and highly reliable connectivity services. The data capability oversees data collection, cleansing, processing, and storage within the end-to-end network, offering data subscription and update services to other planes. The computation capability is tasked with scheduling and optimizing various computational resources within the network to handle computational tasks. The security capability provides inherent security support for the entire network, offering security services to customers while opening up security capabilities to external partners or facilitating crowdsourcing of external security capabilities. The AI capability provides an AI engine that, in conjunction with functional requirements across various domains within the network, offers big data analytics, AI algorithms, model training services, and simulation verification of related solutions. AI capabilities can be centralized or distributed, which can be distributed on network elements and terminals to support real-time AI applications, and can also be concentrated on the cloud to implement complex algorithms with high computing requirements. Additionally, the network’s AI capabilities can be opened to the outside world, while external AI capabilities can be introduced into the network, enabling the crowdsourcing of AI capabilities. The management & orchestration capability is responsible for scheduling various capabilities within the network to achieve flexible organization, fulfilling the orchestration and management of different resources and functions to meet users’ personalized service demands.
@@
Figure 14: Architecture diagram of NET4AI. The 6G network physically comprises three entity networks: the cloud CN (including centralized data/computation centers), the edge-side access network (including 3GPP and non-3GPP connection methods), and the terminal-side access network (including wired and wireless connection methods), as depicted in the lower half of the figure. NET4AI represents the mapping and supporting role of communication connectivity, data, computing, security, and management orchestration within AI, among the six capabilities and services (i.e., communication connectivity, data, computation, security, AI, and management & orchestration) abstracted from the 6G network’s functional services, as indicated by the arrows in the upper half of the figure.